OpenStudy (taishmara_016):

A motor is used to lift a person sitting on a platform that , combined weights 650 N . The platform is lifted 25.0 m strait up in 18.0 seconds. What power is developed by the motor in watts? I need to know the -givens -unknowns -equation -substitute -solve

OpenStudy (irishboy123):

\(P = \dfrac{\Delta W}{\Delta t} = \dfrac{F \cdot \Delta s}{\Delta t} = \dfrac{mg \cdot \Delta h}{\Delta t}\) _____________________________________________________________________________ "Die Mathematik ist die Königin der Wissenschaften und die Zahlentheorie ist die Königin der Mathematik."

likeabossssssss (likeabossssssss):

^^